# Planar Collisionless Shock Simulations with Semi-Implicit Particle-in-Cell Model FLEKS

### v1

@@ 3 Discussion
In-situ measurements of Earth’s and Saturn’s bow shocks (summarized in J. C. Raymond et al. [2023], Figure 1) reveal that the downstream electron-to-ion temperature ratio $T_{e}/T_{i}$ decreases from 1 at low Alfvén Mach numbers $\text{M}_{\text{A}}$ to around 0.1-0.3 at $\text{M}_{\text{A}}\sim 10$. This trend is qualitatively consistent with a statistical study of interplanetary shocks by [L. B. Wilson et al., 2020], which showed a large scatter but an average decrease in $T_{e}/T_{i}$ from 1 at low Mach numbers to $\sim 0.05$ at $\text{M}_{\text{A}}\sim 10$, with little dependence on upstream $\beta$ or shock obliquity $\theta_{Bn}$. For $\text{M}_{\text{A}}=7$, as in our simulations, $T_{e}/T_{i}$ falls within the range of 0.2-0.4, aligning with predictions from several perpendicular shock simulations summarized in J. C. Raymond et al. [2023] (Figure 2). For quasi-perpendicular shocks, we find comparable $T_{e}/T_{i}\simeq 0.28$ in 1D (Figure 1c) and $T_{e}/T_{i}\simeq 0.26$ in 2D with in-plane magnetic field geometry (Figure 5f). In our quasi-parallel shock simulations, while both electron and ion temperatures increase, the $T_{e}/T_{i}$ ratios (0.42 in 1D and 0.33 in 2D) are higher than in the corresponding quasi-perpendicular cases. This indicates preferential heating in quasi-parallel shocks and hints that 2D simulations predict higher $T_{e}/T_{i}$ ratios compared to 1D simulations.
@@
The presence of an upstream-moving diffuse ion population extending to the inflow boundary raises questions about the ultimate fate of these reflected ions. While they energize waves and may be scattered back downstream, our simulations consistently show some ions reaching the boundary. To maintain stability, we enforce a fixed Maxwellian distribution in the boundary ghost cells, effectively removing these ions. This artificial boundary condition, while not causing significant issues in our local planar shock simulations, represents an important consideration for future work, particularly in the context of global MHD-AEPIC coupled simulations. Investigating the transmission of waves and particles through the MHD-AEPIC boundary and its influence on the coupled system will be crucial for accurately capturing the global dynamics of the magnetosphere. For a detailed discussion of the current MHD-AEPIC coupling scheme in our model and related fast and whistler wave tests, interested readers are referred to L. K. S. Daldorff et al. [2014]; Y. Shou et al. [2021]; Y. Chen et al. [2023].
@@ 4 Conclusion and Future Work
This study presents new applications of the improved semi-implicit energy-conserving particle-in-cell (PIC) model FLEKS, focusing on collisionless shock simulations. The refined semi-implicit energy-conserving algorithm in FLEKS proved robust and efficient, enabling stable and accurate shock simulations with grid resolutions at sub-ion scales. We conducted systematic one- and two-dimensional numerical experiments to validate the model under typical terrestrial magnetosphere shock configurations. Our simulations successfully captured the characteristic features of both quasi-perpendicular and quasi-parallel shocks, including shock structures, wave generation, and particle dynamics. Importantly, we highlighted the necessity of at least two spatial dimensions for accurately capturing the full physics of collisionless shocks.
The results of these parameter studies provide valuable guidance for selecting physical and numerical parameters in global magnetosphere simulations. Given current computational capabilities, performing three-dimensional global shock simulations with realistic mass ratios and fully resolved electron physics remains challenging. However, using a reduced mass ratio, even at $m_{i}/m_{e}=25$, can still yield meaningful insights into kinetic processes as revealed from local shock simulations. We anticipate that global magnetosphere simulations incorporating full ion kinetics and appropriately modified electron physics will be instrumental in bridging the gap between large-scale MHD phenomena and small-scale kinetic processes.
@@
Furthermore, to address more realistic shock geometries, large-scale 2D/3D global magnetosphere simulations are required to investigate the dynamics of the curved bow shock and magnetosheath jets. This represents a crucial step beyond planar shock simulations, which can further capture the spatial separation of electron and ion foreshock regions due to the difference of electron and ion masses and time-of-flight effects. Global simulations will enable a more comprehensive understanding of the interplay between shock geometry, particle dynamics, and wave generation in the Earth’s magnetosphere.
We thank Dr. Adnane Osmane, Dr. Yufei Hao, Dr. Xin An, and Dr. Zhongwei Yang for the insightful discussions on the shock wave properties. We gratefully acknowledge NASA Living With A Star (Grant 80NSSC24K0144 and 80NSSC23K1409-LWS) for the financial support. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. {contribution} HZ came up with the initial research concept and was responsible for designing the tests, validating the results, writing and submitting the manuscript. YC obtained the funding and designed the updated algorithm. CD supervised the project and provided ideas for application in global studies. LW conducted comparative simulations to validate the results. YZ also provided the funding and ideas about downstream wave studies. BW provided feedbacks on the modeling results and raised interesting discussions about energy transport. GT offered suggestions on algorithm improvements and insights in interpreting the scalings. All authors edited and contributed to the manuscript.

### v2

@@ 4.5 Electron-to-Ion Temperature Ratio
In-situ measurements of Earth’s and Saturn’s bow shocks (summarized in J. C. Raymond et al. [2023], Figure 1) reveal that the downstream electron-to-ion temperature ratio $T_{e}/T_{i}$ decreases from 1 at low Alfvén Mach numbers $\text{M}_{\text{A}}$ to around 0.1-0.3 at $\text{M}_{\text{A}}\sim 10$. This trend is qualitatively consistent with a statistical study of interplanetary shocks by [L. B. Wilson et al., 2020], which showed a large scatter but an average decrease in $T_{e}/T_{i}$ from 1 at low Mach numbers to $\sim 0.05$ at $\text{M}_{\text{A}}\sim 10$, with little dependence on upstream $\beta$ or shock obliquity $\theta_{Bn}$. For the $\text{M}_{\text{A}}=7$ shocks in our simulations, $T_{e}/T_{i}$ falls within the range of 0.2-0.4, aligning with predictions from several perpendicular shock simulations summarized in J. C. Raymond et al. [2023] (Figure 2). For quasi-perpendicular shocks, we find comparable $T_{e}/T_{i}\simeq 0.28$ in 1D (Figure 1c) and $T_{e}/T_{i}\simeq 0.26$ in 2D with in-plane magnetic field geometry (Figure 5f). In our quasi-parallel shock simulations, while both electron and ion temperatures increase, the $T_{e}/T_{i}$ ratios (0.42 in 1D and 0.33 in 2D) are higher than in the corresponding quasi-perpendicular cases. This indicates preferential heating in quasi-parallel shocks and hints that 2D simulations predict higher $T_{e}/T_{i}$ ratios compared to 1D simulations.
4.6 Numerical Performance and Model Limitations
Our grid resolution study (Figure 4) highlights the stability of the semi-implicit PIC algorithm. The coarsest resolution $\Delta x=4\,d_{e,\text{up}}\simeq 1\,d_{i,\text{up}}$ $(m_{i}/m_{e}=25)$ is close to the ion inertial length. Explicit full-PIC methods generally cannot work at this resolution due to severe finite-grid instability and numerical heating [C. K. Birdsall & A. B. Langdon, 2018]. In contrast, FLEKS produces stable solutions that stay meaningful at sub-ion scales, demonstrating the algorithm’s robustness in overcoming the constraints of explicit PIC methods.
@@ 4.6 Numerical Performance and Model Limitations
A remaining consideration is the presence of an upstream-moving diffuse ion population extending to the inflow boundary in the quasi-parallel runs (Figure 6s-t, Figure 8). While they energize waves and may be scattered back downstream, our simulations consistently show some ions reaching the boundary. To maintain stability, we enforce a fixed Maxwellian distribution in the boundary ghost cells, effectively removing these ions. This artificial boundary condition, while not causing significant issues in our local planar shock simulations, represents an important consideration for future work, particularly in the context of global MHD-AEPIC coupled simulations. Investigating the transmission of waves and particles through the MHD-AEPIC boundary and its influence on the coupled system will be crucial for accurately capturing the global dynamics of the magnetosphere. For a detailed discussion of the current MHD-AEPIC coupling scheme in our model and related fast and whistler wave tests, interested readers are referred to L. K. S. Daldorff et al. [2014]; Y. Shou et al. [2021]; Y. Chen et al. [2023].
@@ 5 Conclusion and Future Work
This study presents new applications of the improved semi-implicit energy-conserving particle-in-cell (PIC) model FLEKS, focusing on collisionless shock simulations. The refined semi-implicit energy-conserving algorithm in FLEKS proved robust and efficient, enabling stable and accurate shock simulations with grid resolutions at sub-ion scales. We conducted systematic one- and two-dimensional numerical experiments to validate the model under typical hypersonic, $\beta\sim 1$ shocks observed at the terrestrial magnetosphere and other astrophysical systems. Our simulations successfully captured the characteristic features of both quasi-perpendicular and quasi-parallel shocks, including shock structures, wave generation, particle dynamics and downstream magnetic reconnection. Importantly, we highlighted the necessity of at least two spatial dimensions for accurately capturing the full physics of collisionless shocks.
The results of these parameter studies provide valuable guidance for selecting physical and numerical parameters in global supersonic, super-Alfvénic magnetosphere simulations. Given current computational constraints, performing three-dimensional global shock simulations with realistic mass ratios and fully resolved electron physics remains challenging. However, using a reduced mass ratio, even at $m_{i}/m_{e}=25$, can still yield meaningful insights into kinetic processes as revealed from the semi-implicit PIC local shock simulations. We anticipate that global magnetosphere simulations incorporating full ion kinetics and appropriately modified electron physics will be instrumental in bridging the gap between large-scale MHD phenomena and small-scale kinetic processes.
@@
Furthermore, to address more realistic bow shock geometries, large-scale 2D/3D coupled global magnetosphere simulations are required to investigate the dynamics of the curved shocks and downstream jets. This represents a crucial step beyond planar shock simulations, which can further capture the spatial separation of electron and ion foreshock regions due to the difference of electron and ion masses and time-of-flight effects. Global simulations will enable a more comprehensive understanding of the interplay between shock geometry, particle dynamics, and wave generation in the planetary magnetospheres.
This work was supported by NASA grants 80NSSC23K1409 and 80NSSC24K0144, DOE grant DE-SC0024639 and the Alfred P. Sloan Research Fellowship. We thank Dr. Adnane Osmane, Dr. Yufei Hao, Dr. Xin An, and Dr. Zhongwei Yang for the insightful discussions on the shock wave properties. We also thank the reviewer for the constructive comments that greatly improve the manuscript. We would like to acknowledge high-performance computing support from the NASA High-End Computing Program through the NASA Advanced Super-computing Division at Ames Research Center, from National Energy Research Scientific Computing Center, a DOE Office of Science user facility, and from the Derecho system (doi:10.5065/qx9a-pg09) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. {contribution}HZ came up with the initial research concept and was responsible for designing the tests, validating the results, writing and submitting the manuscript. YC obtained the funding and designed the updated algorithm. CD supervised the project and provided ideas for application in global studies. LW conducted comparative simulations to validate the results. YZ also provided the funding and ideas about downstream wave studies. BW provided feedbacks on the modeling results and raised interesting discussions about energy transport. GT offered suggestions on algorithm improvements and insights in interpreting the scalings. All authors reviewed and contributed to the manuscript.